# Neural Pushforward Samplers for the Fokker–Planck Equation on Embedded Riemannian Manifolds

### 17 Mar 2026 v1

@@ Abstract
We extend the Weak Adversarial Neural Pushforward Method (WANPF) to the Fokker–Planck equation posed on a compact, smoothly embedded Riemannian manifold $M\hookrightarrow\mathbb{R}^{n}$. The key observation is that the weak formulation of the Fokker–Planck equation, together with the ambient-space representation of the Laplace–Beltrami operator via the tangential projection $P(x)$ and the mean-curvature vector $H(x)$, permits all integrals to be evaluated as expectations over samples lying on $M$, using test functions defined globally on $\mathbb{R}^{n}$. A neural pushforward map $F_{{\boldsymbol{\vartheta}}}$ is constrained to map the support of a base distribution into $M$ at all times through a manifold retraction $\Pi_{M}$, so that probability conservation and manifold membership are enforced by construction. Adversarial ambient plane-wave test functions are chosen, and their Laplace–Beltrami operators are derived in closed form, enabling autodiff-free, mesh-free training. We present both a steady-state and a time-dependent formulation, derive explicit Laplace–Beltrami formulae for the sphere $S^{n-1}$ and the flat torus $\mathbb{T}^{n}$, and demonstrate the method numerically on a double-well steady-state Fokker–Planck equation on $S^{2}$. The trained generator concentrates mass near the two potential minima at $(\pm 1,0,0)$, consistent with the Boltzmann distribution of the underlying Langevin dynamics.

### 18 Mar 2026 v2

@@ Abstract
In this paper, we extend the weak adversarial neural pushforward mapping method to the Fokker–Planck equation on compact embedded Riemannian manifolds. The method represents the solution as a probability distribution via a neural pushforward map that is constrained to the manifold by a retraction layer, enforcing manifold membership and probability conservation by construction. Training is guided by a weak adversarial objective using ambient plane-wave test functions, whose intrinsic differential operators are derived in closed form from the geometry of the embedding, yielding a fully mesh-free and chart-free algorithm. Both steady-state and time-dependent formulations are developed, and numerical results on a double-well problem on the two-sphere demonstrate the capability of the method in capturing multimodal invariant distributions on curved spaces.